# Deep Reinforcement Learning for Robotics: A Survey of Real-World Successes

### v2

@@ 4.3.3 In-hand Manipulation
Humans exhibit many in-hand manipulation behaviors, re-orienting and re-positioning objects to facilitate downstream manipulation. Impressive strides in the development of these capabilities have been made with DRL in recent years, allowing agents to learn such complex in-hand manipulation behaviors with impressive generalization. Several works focused on re-orienting single objects to target configurations [138, 139], employing pose estimation modules trained in simulation. Nagabandi et al. [140] similarly demonstrated rotating Baoding balls with model-based RL. While showing impressive dexterity, these works focus on manipulating known objects (e.g., a given cube) with low-dimensional observations. Recent methods leveraging vision and tactile data have demonstrated rotating arbitrary objects about arbitrary axes [141], even against gravity [142]. These approaches employ extensive domain randomization and typically leverage privileged information (e.g., object shape information, dynamic properties) and dense rewards when training in simulation. An open challenge is integrating these in-hand manipulation skills with other manipulation abilities (e.g., tool use), which require re-orientation to a target configuration suitable for a downstream task.
@@ 4.4.3 Long-horizon Interactive Tasks
WBC 152,153,154,155
Short-Horizon 158,159,160,161,162,163,164,165,166,167,168,169
Long-Horizon 157,170,171
@@ 4.5.2 Non-collaborative pHRI
Collaborative pHRI 173, 172, 174, 180
Non-collaborative pHRI 175, 176, 177, 178, 179
Shared Autonomy 181, 182, 183
@@ 4.6.1 Multi-Agent Collision Avoidance
Multi-Robot Collision Avoidance 184, 185, 187, 188, 189
Multi-Robot Loco-Manipulation 190
Robot Soccer 191
@@ Appendix B Additional Tables
Manipulation 113, 122, 127, 131, 138, 139, 140, 141, 142 54, 110, 111, 112, 114, 115, 116, 117, 118, 119, 120, 121, 124, 125, 126, 128, 129, 130, 132, 133, 134, 137, 143, 144, 145 108, 109, 123, 135, 136
MoMa 153, 154, 155, 163, 167, 169, 158, 162, 159, 166 152, 164, 160, 161, 171, 157, 170, 165 168
HRI 175, 176, 177, 178, 179, 180 173, 174, 181, 182, 183 172
Multi-Robot Interaction 184, 185, 187, 188, 190, 191 189
@@
Manipulation 54, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 124, 125, 128, 133, 134, 135, 136, 137, 141, 142, 143, 144, 145 122, 123, 126, 127, 129, 130, 131, 132, 138, 139, 140 54, 108, 110, 111, 112, 114, 115, 117, 118, 122, 124, 128, 129, 134 109, 113, 116, 118, 119, 120, 121, 123, 125, 126, 127, 130, 131, 132, 133, 135, 136, 137, 138, 139, 140 141, 142, 143 144, 145
MoMa 163, 167, 164, 160, 168, 165, 159, 171, 157, 170 153, 154, 152, 155, 169, 158, 162, 161, 166 168, 171, 170 153, 154, 152, 155, 163, 167, 169, 158, 162, 164, 160, 161, 165, 159, 166, 157
HRI 173, 174, 177, 178, 181 172, 175, 176, 179, 180, 182, 183 172 173, 174, 175, 176, 177, 178, 179, 180, 181, 182, 183
Multi-Robot Interaction 184, 185, 187, 188, 189, 190, 191 184, 185 187, 188, 189, 190, 191
@@
Manipulation 108, 111, 123, 130, 133, 134, 135, 137, 138, 139, 141, 142, 143, 144, 145 116, 131 54, 109, 110, 112, 113, 114, 115, 117, 118, 119, 120, 121, 122, 124, 125, 126, 127, 128, 129, 132, 136, 140
MoMa 153, 154, 152, 155, 163, 167, 169, 162, 164, 161, 165, 159, 166, 171, 157 158, 170 160, 168
HRI 172, 173, 174, 175, 176, 177, 179 182 178, 181, 180
Multi-Robot Interaction 184, 185, 187, 188, 189, 190, 191
@@
Manipulation 54, 113, 118, 140 108, 109, 110, 111, 112, 114, 115, 116, 117, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 141, 142, 143, 144, 145 54, 108, 109, 110, 111, 114, 115, 116, 118, 119, 123, 125, 130, 131, 133, 135, 137, 138, 139, 140, 141, 142, 143, 144, 145 112, 113, 117, 120, 121, 122, 124, 126, 127, 128, 129, 132, 134, 136
MoMa 153, 154, 152, 155, 163, 167, 169, 158, 162, 164, 160, 161, 168, 165, 159, 166, 171, 157, 170 153, 154, 152, 163, 167, 169, 158, 162, 164, 160, 161, 165, 159, 166, 171 155, 168, 157, 170
HRI 173, 174, 181 172, 175, 176, 177, 178, 179, 180, 182, 183 173, 174, 177, 179, 180, 182, 183 172, 175, 176, 178, 181
Multi-Robot Interaction 184, 185, 187, 188, 189, 190, 191 187, 188, 190 184, 185, 189, 191
@@
Manipulation 118, 140 108, 115, 121, 122, 124, 129 54, 109, 110, 111, 112, 114, 116, 117, 119, 120, 123, 125, 126, 127, 128, 132, 133, 134, 135, 136, 137, 143, 144 113, 130, 131, 138, 139, 141, 142, 145
MoMa 158, 164, 160, 161, 171, 170 153, 154, 152, 155, 163, 167, 169, 158, 162, 168, 165, 159, 166, 157
HRI 181 178 172, 173, 174, 178, 179, 180, 182 175, 176, 177, 183
Multi-Robot Interaction 191 184, 185, 187, 188, 189, 190
@@
Manipulation 123, 129, 131, 132, 138, 139, 140, 143 54, 108, 109, 110, 111, 112, 113, 114, 115, 117, 118 119, 120, 121, 122, 125, 126, 127, 128, 133, 134, 135, 136, 137, 142, 144 54, 118, 130 116, 124, 141, 145
MoMa 153, 154, 152, 155, 169, 158, 162, 161, 166 163, 167, 164, 160, 168, 165, 159, 171, 157, 170 167, 165, 170
HRI 175, 179, 180, 182, 183 173, 174, 177, 178, 181 172, 176 181
Multi-Robot Interaction 184, 187, 190, 191 185, 188, 189

### v3

@@ 4.3.3 In-hand Manipulation
Humans exhibit many in-hand manipulation behaviors, re-orienting and re-positioning objects to facilitate downstream manipulation. Impressive strides in the development of these capabilities have been made with DRL in recent years, allowing agents to learn such complex in-hand manipulation behaviors with impressive generalization. Several works focused on re-orienting single objects to target configurations [138, 139], employing pose estimation modules trained in simulation or using proprioception alone [143]. Nagabandi et al. [140] similarly demonstrated rotating Baoding balls with model-based RL. While showing impressive dexterity, these works focus on manipulating known objects (e.g., a given cube) with low-dimensional observations. Recent methods leveraging vision and/or tactile data have demonstrated rotating arbitrary objects about arbitrary axes [141], even against gravity [142, 144]. These approaches employ extensive domain randomization and typically leverage privileged information (e.g., object shape information, dynamic properties) and dense rewards when training in simulation. An open challenge is integrating these in-hand manipulation skills with other manipulation abilities (e.g., tool use), which require re-orientation to a target configuration suitable for a downstream task.
@@ 4.4.3 Long-horizon Interactive Tasks
WBC 154,155,156,157
Short-Horizon 160,161,162,163,164,165,166,167,168,169,170,171
Long-Horizon 159,172,173
@@ 4.5.2 Non-collaborative pHRI
Collaborative pHRI 175, 174, 176, 182
Non-collaborative pHRI 177, 178, 179, 180, 181
Shared Autonomy 183, 184, 185
@@ 4.6.1 Multi-Agent Collision Avoidance
Multi-Robot Collision Avoidance 186, 187, 189, 190, 191
Multi-Robot Loco-Manipulation 192
Robot Soccer 193
@@ Appendix B Additional Tables
Manipulation 113, 122, 127, 131, 138, 139, 140, 141, 142, 143, 144 54, 110, 111, 112, 114, 115, 116, 117, 118, 119, 120, 121, 124, 125, 126, 128, 129, 130, 132, 133, 134, 137, 145, 146, 147 108, 109, 123, 135, 136
MoMa 155, 156, 157, 165, 169, 171, 160, 164, 161, 168 154, 166, 162, 163, 173, 159, 172, 167 170
HRI 177, 178, 179, 180, 181, 182 175, 176, 183, 184, 185 174
Multi-Robot Interaction 186, 187, 189, 190, 192, 193 191
@@
Manipulation 54, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 124, 125, 128, 133, 134, 135, 136, 137, 141, 142, 145, 146, 147 122, 123, 126, 127, 129, 130, 131, 132, 138, 139, 140, 143, 144 54, 108, 110, 111, 112, 114, 115, 117, 118, 122, 124, 128, 129, 134 109, 113, 116, 118, 119, 120, 121, 123, 125, 126, 127, 130, 131, 132, 133, 135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145 146, 147
MoMa 165, 169, 166, 162, 170, 167, 161, 173, 159, 172 155, 156, 154, 157, 171, 160, 164, 163, 168 170, 173, 172 155, 156, 154, 157, 165, 169, 171, 160, 164, 166, 162, 163, 167, 161, 168, 159
HRI 175, 176, 179, 180, 183 174, 177, 178, 181, 182, 184, 185 174 175, 176, 177, 178, 179, 180, 181, 182, 183, 184, 185
Multi-Robot Interaction 186, 187, 189, 190, 191, 192, 193 186, 187 189, 190, 191, 192, 193
@@
Manipulation 108, 111, 123, 130, 133, 134, 135, 137, 138, 139, 141, 142, 143, 144, 145, 146, 147 116, 131 54, 109, 110, 112, 113, 114, 115, 117, 118, 119, 120, 121, 122, 124, 125, 126, 127, 128, 129, 132, 136, 140
MoMa 155, 156, 154, 157, 165, 169, 171, 164, 166, 163, 167, 161, 168, 173, 159 160, 172 162, 170
HRI 174, 175, 176, 177, 178, 179, 181 184 180, 183, 182
Multi-Robot Interaction 186, 187, 189, 190, 191, 192, 193
@@
Manipulation 54, 113, 118, 140, 144 108, 109, 110, 111, 112, 114, 115, 116, 117, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 141, 142, 143, 145, 146, 147 54, 108, 109, 110, 111, 114, 115, 116, 118, 119, 123, 125, 130, 131, 133, 135, 137, 138, 139, 140, 141, 142, 143, 144, 145, 146, 147 112, 113, 117, 120, 121, 122, 124, 126, 127, 128, 129, 132, 134, 136
MoMa 155, 156, 154, 157, 165, 169, 171, 160, 164, 166, 162, 163, 170, 167, 161, 168, 173, 159, 172 155, 156, 154, 165, 169, 171, 160, 164, 166, 162, 163, 167, 161, 168, 173 157, 170, 159, 172
HRI 175, 176, 183 174, 177, 178, 179, 180, 181, 182, 184, 185 175, 176, 179, 181, 182, 184, 185 174, 177, 178, 180, 183
Multi-Robot Interaction 186, 187, 189, 190, 191, 192, 193 189, 190, 192 186, 187, 191, 193
@@
Manipulation 118, 140 108, 115, 121, 122, 124, 129 54, 109, 110, 111, 112, 114, 116, 117, 119, 120, 123, 125, 126, 127, 128, 132, 133, 134, 135, 136, 137, 143, 145, 146 113, 130, 131, 138, 139, 141, 142, 144, 147
MoMa 160, 166, 162, 163, 173, 172 155, 156, 154, 157, 165, 169, 171, 160, 164, 170, 167, 161, 168, 159
HRI 183 180 174, 175, 176, 180, 181, 182, 184 177, 178, 179, 185
Multi-Robot Interaction 193 186, 187, 189, 190, 191, 192
@@
Manipulation 123, 129, 131, 132, 138, 139, 140, 143, 144, 145 54, 108, 109, 110, 111, 112, 113, 114, 115, 117, 118, 119, 120, 121, 122, 125, 126, 127, 128, 133, 134, 135, 136, 137, 142, 146 54, 118, 130, 144 116, 124, 141, 147
MoMa 155, 156, 154, 157, 171, 160, 164, 163, 168 165, 169, 166, 162, 170, 167, 161, 173, 159, 172 169, 167, 172
HRI 177, 181, 182, 184, 185 175, 176, 179, 180, 183 174, 178 183
Multi-Robot Interaction 186, 189, 192, 193 187, 190, 191